solve the equation for the interval [0,2pi)

sin2X + sin X = 0
1
Expert's answer
2012-12-19T11:06:34-0500
sin2X+ sin X = 0
2sinX cosX + sin X = 0
sin X(2cos X +1 ) = 0
several solutions
sin X = 0
X = 0, Pi
2cos X = -1
cos X = -1/2
X = 2/3 Pi, 7/6Pi